OpenStudy (anonymous):

what is the extraneous solution of -x=2x+15

OpenStudy (solomonzelman):

extraneous solution is a solution that works lets solve for x -x=2x+15 -3x=15 x=-5 Does that work? -(-5) = 2 (-5) + 15 5= -10 + 15 So what's the extraneous solution ?

myininaya (myininaya):

extraneous solution is a solution that works for the simplified equation that doesn't work for the original equation.

OpenStudy (solomonzelman):

I didn't know that this is the definition, tnx for telling me myinimaya. (didn't want to disturb you by tagging)

myininaya (myininaya):

lol It would have been fine.

OpenStudy (solomonzelman):

Alright, ty either way :)